# Can Code-Switched Texts Activate a Knowledge Switch [Uncaptioned image] in LLMs? A Case Study on English-Korean Code-Switching

### v2

@@ Abstract
Seoyeon Kim1    Huiseo Kim1    Chanjun Park2    Jinyoung Yeo1    Dongha Lee1thanks:   Corresponding author 1Yonsei University   2Upstage {emseoyk, pororokhs, jinyeo, donalee}@yonsei.ac.kr chanjun.park@upstage.ai
@@ 1 Introduction
In human multilingual societies, code-switching (CS), or the practice of alternating between two or more languages within an utterance, is used to fill in lack of language proficiency, to emphasize certain emotions or points, or for group identity (Heredia and Altarriba, 2001). Moreover, code-switching functions as an effective tool to embed cultural meanings. Expressing certain concepts in original language can convey subtle cultural and linguistic nuances that can be lost in translation, and knowledge related to certain language are more likely to be more memorized in its own language. As shown in Figure 1, when a human English-Korean bilingual is given a question that is closely related to Korean culture, a question in English and Korean code-switching is more capable of recalling knowledge about “몽유도원도” [^1] A landscape painting by An Gyeon in the early Joseon Dynasty requested by Prince Anpyeong, after his dream about Shangri-la. The painting is drawn on silk with ink., because the concept is more familiar in Korean than in English.
@@
Meanwhile, a crucial challenge when it comes to code-switching is the data scarcity. There is a limited number of CS datasets, let alone culture-focused data (Doğruöz et al., 2021). Since CS often happens in conversations, data are not easily available and the quality is not ensured. To address the shortage of data, efforts have been made to synthetically generate code-switching corpus based on linguistic theories (Pratapa et al., 2018; Rizvi et al., 2021; Salaam et al., 2022). However, these works rely on syntactic parsers and part-of-speech taggers that support limited languages, and the quality of text are highly dependent on the performances of those tools. Therefore, we first construct EnKoQA, a synthetic English-Korean code-switching dataset to explore the potential of CS in low-resource language task. Following Matrix Language Frame Model (Myers-Scotton, 1997), we synthesize Korean QA datasets (Kim et al., 2024b; Son et al., 2024) that encompass various aspects of Korea into English-Korean code-switched questions.
@@ 2.3 Code-Switching for Language Models
Previous works introduce benchmarks for evaluating code-switching ability of multilingual language models across multiple tasks (Aguilar et al., 2020; Khanuja et al., 2020). More recent works focus on the capability of LLMs in code-switching. Zhang et al. (2023a) discover performance of multilingual LLMs in various code-switching tasks, including sentiment analysis and language identification. Yong et al. (2023) explore prompting multilingual LLMs to generate code-mixed data. Shankar et al. (2024) introduce a prompting technique called in-context mixing for effective in-context learning in LLMs. Although these benchmarks encompass a variety of tasks, the analysis of LLMs’ code-switching capabilities in terms of knowledge retrieval and utilization has not yet been investigated.
@@ 2.4 Code-Switched Data Synthesis
Data synthesis for code-switching has been approached in various ways. Several studies utilize parsers and neural models to synthesize code-switched text based on EC theory (Pratapa et al., 2018; Rizvi et al., 2021). Similarly, Salaam et al. (2022) extract phrases from source language and reintegrate them into target language. In recent efforts to address data scarcity in low-resource settings, LLMs have been employed to generate synthetic data (Li et al., 2023). However, using LLMs specifically for synthesizing code-switched data remains unexplored.
@@ 3.2 Experimental Settings
To compare performances of LLMs in various language settings, we evaluate on CS, English, and translated Korean (KOt) questions. Korean translation baseline simulates cases where machine translation is adopted to convert task data from English into the target language in low-resource language tasks. We back-translate English translation text to Korean using gpt-3.5-turbo. Prompts that we used for inference are provided in Table 13. We also conduct experiments on the original Korean questions, but do not consider it as major baseline, because we aim to examine the effect of code-switching compared to dominant language, rather than demonstrate the performance of low-resource language. Please refer to Appendix B.3 for further discussions.
@@ 6 Analysis on Knowledge Leveraging
We demonstrate that qualified knowledge identification is prerequisite for knowledge activation of CS. The win ratio of History knowledge by GPT-3.5 was relatively poor compared to others (Figure 4), leading GPT-3.5 to be the only model that did not benefit from CS. Similarly in Law, Figure 2 and 3 show that helpfulness and pairwise scores for knowledge by Gemma2 9B and 27B are lower than others, which are responsible for their suboptimal performance in CS. We provide further qualitative case study in Appendix D.4.
@@
Although we informed the models that the answer is in one of the choices, we notice that the majority of incorrect responses were “None of the above”. The errors may derive from either hallucinated knowledge or failing to follow instructions faithfully. Therefore, we provide additional analysis on erroneous outputs in Table 8. We report the results in the format of number of errors that derived from knowledge hallucination / total number of None errors. Errors that are not from hallucination are caused by poor instruction-following. Overall, we observe that performance on English questions results in more errors compared to CS across all LLMs, and most of them were hallucination errors. This indicates that models hallucinate much frequently when English questions are given, again highlighting the effectiveness of CS over English setting. It is also worth noting that Gemma2 families hallucinate largely on History and General, supporting our finding in Figure 2 and Figure 5 which respectively illustrates poor performance on human evaluation and QA accuracy.
7 Conclusion
@@
We explore the efficacy of code-switching in activating language-specific knowledge embedded in LLMs. Utilizing two Korean-centric QA datasets, we synthesize EnKoQA, a qualified English-Korean code-switching QA dataset and conduct experiments on various multilingual LLMs. Our analyses demonstrate that LLMs can simulate a similar code-switching effect with human communications of facilitating low-resource knowledge within LLMs, particularly in language-specific domains. Regarding this finding, we suggest that code-switching can be an effective strategy for solving low-resource language tasks. Also, augmenting low-resource datasets into code-switching text can amplify resource and mitigate data scarcity challenge. We hope our work can motivate NLP community to explore more potential of code-switching for developing robust multilingual LLMs.
Limitations and Future Work
@@
In this work, we focus on code-switching between English and Korean, specifically limiting the scope to Korea-specific knowledge. However, it is important to note that this study serves as a single case focused on the Korean context and leaves room for expanding the scope of code-switching to other cultures and languages. For future research, we aim to investigate whether the knowledge activation effect also occurs in other language settings.
@@
In the future, we aim to investigate more potential of code-switching in diverse aspects, including instruction-tuning of LLMs to users effectively using code-switching for multilingual tasks. As we have demonstrated synthesizing monolingual datasets into code-switching text, we hope our work can inspire NLP community to explore the capability of code-switching in enhancing and utilizing multilingual LLMs.
@@ A.5 Annotation Details
Regarding inter-annotator agreement (IAA), although we did not compute a formal IAA score, significant effort was devoted to ensuring high annotation quality through extensive discussion and collaboration among annotators. In specific, the annotation process involved annotators who are fluent in both English and Korean are assigned each portion of the dataset to select a candidate for code-switched question. Following this initial annotation, the annotators cross-checked each other’s work to identify any discrepancies. If a selected candidate appeared to be incorrect or suboptimal, the annotators engaged in thorough discussions until they reached an agreement on the most appropriate candidate. This iterative and collaborative process was integral to constructing a high-quality dataset.
@@ A.6 Dataset Size and Quality
KO 다음 글의 (가)에 대한 (나)의 상대적 특성으로 옳은 것은? (단, (가), (나)는 각각 겨울과 여름 중 하나임.) 우리나라는 더위와 추위에 대비하여 대청마루와 온돌 같은 전통 가옥 시설이 발달하였다. 대청마루는 바람을 잘 통하게 하여 (가) 을 시원하게 지낼 수 있도록 설치되었다. 온돌은 아궁이의 열을 방으로 전달하여 (나) 을 따뜻하게 지낼 수 있도록 설치되었다. 대청마루 는 중부와 남부 지역에 발달한 한편, 온돌은 대부분의 지역에 발달하였다. (1) 평균 상대 습도가 높다. (2) 정오의 태양 고도가 높다. (3) 한파의 발생 일수가 많다. (4) 대류성 강수가 자주 발생한다. (5) 열대 저기압의 통과 횟수가 많다.
EN What is the correct relative characteristic of (나) in relation to (가) in the following passage? (Note that (가) and (나) refer to either winter or summer.) In Korea, traditional house facilities such as daecheongmaru and ondol have developed to cope with heat and cold. Daecheongmaru is designed to allow good ventilation to keep (가) cool. Ondol transfers heat from the kitchen stove to the room to keep (나) warm. While daecheongmaru is developed in the central and southern regions, ondol is developed in most areas. (1) The average relative humidity is high. (2) The midday sun’s altitude is high. (3) There are many days of occurrence of cold waves. (4) Heavy rainfall often occurs in Daeryuseong. (5) There are many occurrences of passage of tropical cyclones.
CS What is the correct relative characteristic of (나) in relation to (가) in the following passage? (Note that (가) and (나) refer to either winter or summer.) In 한국, 전통 가옥 시설 such as 대청마루 and 온돌 have developed to cope with heat and cold. 대청마루 is designed to allow good ventilation to keep (가) cool. 온돌 transfers heat from the kitchen stove to the room to keep (나) warm. While 대청마루 is developed in the 중부 and 남부 지역, 온돌 is developed in most areas. (1) The average 상대 습도 is high. (2) The 정오의 태양 고도 is high. (3) There are many days of occurrence of 한파. (4) 대류성 강수 often occurs. (5) There are many occurrences of passage of 열대 저기압.
@@ C.2 Human Evaluator Qualifications
For knowledge identification evaluation, collecting qualified bilingual evaluators was not easy due to the inherent challenge in code-switching research of necessitating fluent bilinguals as evaluators. Our dataset is composed of questions from Korean proficiency tests for foreigners and the Korean College Scholastic Ability Test. Thus, it is designed at a level that would not be challenging for evaluators whom were born and raised in Korea, received a Korean public education, and graduated prestigious universities. We managed to collect four Korean graduate school students as our evaluators, all of whom are native Korean with sufficient understanding of Korean culture. Also, they possess qualified English exam scores, indicating that they have no problem in understanding Korean-English code-switched texts. To mitigate the shortage of labor force, we designed the evaluation criteria objectively, allowing for an assessment that is not subjective and has clear correct answers. Specifically, we evaluate knowledge identification based on two criteria: faithfulness and helpfulness. Faithfulness evaluates the factualness of the knowledge, so the evaluators are required to use their background knowledge as well as searching from faithful sources where gold knowledge exists. To evaluate helpfulness, evaluators are given a gold answer to the question and determine whether the knowledge is helpful for finding the answer, using their logical reasoning.
@@ Appendix D Observations
Model History Tradition General Law
GPT-4o 0.41 0.64 0.62 0.62
Solar 0.26 -0.09 0.38 0.02
Gemma2 27B 0.25 0.52 0.17 0.34
Gemma2 9B 0.20 -0.07 0.05 0.24
@@ D.3 Error Analysis
D.4 Case Study
We examine a sample case to compare the capability of code-switching and English on knowledge activation. Table 10 shows the knowledge and answer generated by Solar in Tradition. The question asks about “정월대보름”, a Korean traditional holiday that celebrates the first full moon of lunar new year. We observe that CS question preserves unique terms such as “정월대보름” and ”귀밝이술” in Korean; this helps the model to successfully activate faithful knowledge, consequently leading to the correct answer. However, in the case of English, not only are these cultural nuances lost in English question, but the model misunderstood the question to asking about “Dan-o”, another Korean traditional holiday. Solar lacks in knowledge about “정월대보름” in English, or fails to activate encoded knowledge with its English translation.

### v3

@@ Abstract
Seoyeon Kim1    Huiseo Kim1    Chanjun Park2    Jinyoung Yeo1    Dongha Lee1thanks:   Corresponding author 1Yonsei University   2Soongsil University {emseoyk, pororokhs, jinyeo, donalee}@yonsei.ac.kr chanjun.park@ssu.ac.kr
@@ 1 Introduction
In human multilingual societies, code-switching (CS), or the practice of alternating between two or more languages within an utterance, is used to fill in lack of language proficiency, to emphasize certain emotions or points, or for group identity (Heredia and Altarriba, 2001). Moreover, code-switching functions as an effective tool to embed cultural meanings. Expressing certain concepts in original language can convey subtle cultural and linguistic nuances that can be lost in translation, and knowledge related to certain language are more likely to be more memorized in its own language. As shown in Figure 1, when a human English-Korean bilingual is given a question that is closely related to Korean culture, a question in English and Korean code-switching is more capable of recalling knowledge about ‘몽유도원도’ [^1] A landscape painting by An Gyeon, commissioned by Prince Anpyeong in the early Joseon Dynasty following his dream of Shangri-la., because the concept is more familiar in Korean than in English.
@@
Meanwhile, a crucial challenge when it comes to code-switching is the data scarcity. There is a limited number of CS datasets, let alone culture-focused data (Doğruöz et al., 2021). Since CS often happens in conversations, data are not easily available and the quality is not ensured. To address the shortage of data, efforts have been made to synthetically generate code-switching corpus based on linguistic theories (Pratapa et al., 2018; Rizvi et al., 2021; Salaam et al., 2022). However, these works rely on syntactic parsers and part-of-speech taggers that support limited languages, and the quality of text are highly dependent on the performances of those tools. Therefore, we first construct EnKoQA, a synthetic English-Korean code-switching dataset to explore the potential of CS in low-resource language task. [^2] We further discuss the resource scarcity of Korean language in Appendix A.8. Following Matrix Language Frame Model (Myers-Scotton, 1997), we synthesize Korean QA datasets (Kim et al., 2024b; Son et al., 2024) that encompass various aspects of Korea into English-Korean code-switched questions.
@@ 2.3 Code-Switching for Language Models
Previous works introduce benchmarks for evaluating code-switching ability of multilingual language models across multiple tasks (Aguilar et al., 2020; Khanuja et al., 2020). More recent works focus on the capability of LLMs in code-switching. Zhang et al. (2023a) discover performance of multilingual LLMs in various code-switching tasks, including sentiment analysis and language identification. Yong et al. (2023) explore prompting multilingual LLMs to generate code-mixed data. Shankar et al. (2024) introduce a prompting technique called in-context mixing for effective in-context learning in LLMs. Wang et al. (2025) explores the impact of code-switching on cross-lingual transfer during pre-training. Although these benchmarks encompass a variety of tasks, the analysis of LLMs’ code-switching capabilities in terms of knowledge retrieval and utilization has not yet been investigated.
@@ 2.4 Code-Switched Data Synthesis
Data synthesis for code-switching has been approached in various ways. Several studies utilize parsers and neural models to synthesize code-switched text based on EC theory (Pratapa et al., 2018; Rizvi et al., 2021). Similarly, Salaam et al. (2022) extract phrases from source language and reintegrate them into target language. Another line of works synthesize code-switching on token-level (Li et al., 2024; Wang et al., 2025). In recent efforts to address data scarcity in low-resource settings, LLMs have been employed to generate synthetic data (Li et al., 2023). However, using LLMs specifically for synthesizing code-switched data remains unexplored.
@@ 3.2 Experimental Settings
To compare performances of LLMs in various language settings, we evaluate on CS, English, and translated Korean (KOt) questions. Korean translation baseline simulates more practical scenarios of code-switching where machine translation is adopted to convert task data from English into the target language in low-resource language tasks. This baseline accounts for the possibility of errors that can arise during machine translation, and to compare whether code-switching would produce more robust results in comparison. To create KOt, we back-translate English translation text to Korean using gpt-3.5-turbo. Prompts that we used for inference are provided in Table 13. We also conduct experiments on the original Korean questions, but do not consider it as major baseline, because we aim to examine the effect of code-switching compared to dominant language, rather than demonstrate the performance of low-resource language. Please refer to Appendix B.3 for further discussions.
@@ 5.1 Human Evaluation
Model History Tradition General Law
GPT-4o 0.41 0.64 0.62 0.62
Solar 0.26 -0.09 0.38 0.02
Gemma2 27B 0.25 0.52 0.17 0.34
Gemma2 9B 0.20 -0.07 0.05 0.24
@@ 6 Analysis on Knowledge Leveraging
We demonstrate that qualified knowledge identification is prerequisite for knowledge activation of CS. The win ratio of History knowledge by GPT-3.5 was relatively poor compared to others (Figure 4), leading GPT-3.5 to be the only model that did not benefit from CS. Similarly in Law, Figure 2 and 3 show that helpfulness and pairwise scores for knowledge by Gemma2 9B and 27B are lower than others, which are responsible for their suboptimal performance in CS.
@@
Although we informed the models that the answer is in one of the choices, we notice that the majority of incorrect responses were “None of the above”. The errors may derive from either hallucinated knowledge or failing to follow instructions faithfully. Therefore, we provide additional analysis on erroneous outputs in Table 8. We report the results in the format of # of errors that derived from knowledge hallucination / total # of None errors. Errors that are not from hallucination are caused by poor instruction-following. Overall, we observe that answering English questions results in more errors compared to CS across all LLMs, and most of them were hallucinations. This indicates that models hallucinate much frequently when English questions are given, again highlighting the effectiveness of CS over English. It is also worth noting that Gemma2 families hallucinate largely on History and General, supporting our finding in Figure 2 and 5 which respectively illustrates poor performance on human evaluation and QA accuracy.
Case Study.
We examine a sample case to compare the capability of code-switching and English on knowledge activation. Table 10 shows the knowledge and answer generated by Solar in Tradition. The question asks about ‘정월대보름 (Jeong-wol Dae-bo-reum)’, a Korean traditional holiday that celebrates the first full moon of lunar new year. We observe that CS question preserves unique terms such as ‘정월대보름’ and ‘귀밝이술 (Gwi-bal-ki-sul)’ in Korean; this helps the model to successfully activate faithful knowledge, consequently leading to the correct answer. However, in the case of English, not only are these cultural nuances lost in English question, but the model misunderstood the question to asking about ‘단오 (Dan-o)’, another Korean traditional holiday. Solar lacks in knowledge about ‘정월대보름’ in English, or fails to activate encoded knowledge with its English translation.
7 Conclusion and Future Work
@@
We explore the efficacy of code-switching in activating language-specific knowledge embedded in LLMs. Utilizing two Korean-centric QA datasets, we synthesize EnKoQA, a qualified English-Korean code-switching QA dataset and conduct experiments on various multilingual LLMs. Our analyses demonstrate that LLMs can simulate a similar code-switching effect with human communications of facilitating low-resource knowledge within LLMs, particularly in language-specific domains. Regarding this finding, we suggest that code-switching can be an effective strategy for solving low-resource language tasks. Also, augmenting low-resource datasets into code-switching text can amplify resource and mitigate data scarcity challenge.
Future directions of this work involve a more comprehensive investigation into the potential of code-switching across different aspects of multilingual modeling, including its role in pretraining and instruction-tuning, as well extending our work to a wider range of languages. By first addressing the relatively understudied Korean-English pair, our research contributes to a relatively unexplored area and serves as a stepping stone for subsequent studies on code-switching. In particular, our data synthesis approach, grounded in the universally applicable MLF theory, extends beyond Korean and is generalizable to all languages. Through our work, we encourage the NLP community to further explore the capacity of code-switching to effectively enhance and harness the potential of multilingual LLMs.
Limitations
@@
In this work, we focus on code-switching between English and Korean, specifically limiting the scope to Korea-specific knowledge. However, it is important to note that this study serves as a starting point focused on the Korean context and leaves room for expanding the scope of code-switching to other cultures and languages. For future research, we aim to investigate whether the knowledge activation effect also occurs in other language settings.
@@ A.5 Annotation Details
Regarding inter-annotator agreement (IAA), although we did not compute a formal IAA score, significant effort was devoted to ensuring high annotation quality through extensive discussion and collaboration among annotators. The task for human annotators in our datasets was not labeling answers, but verifying the quality of the dataset. IAA scores are for quantifying annotator agreement and further improving the dataset. In that sense, our qualitative human verification comparably guarantees both annotator consensus and dataset quality. We refined the dataset until it reached perfect agreement between the annotators, which is comparable to 1 in IAA score. In specific, the annotators who are fluent in both English and Korean are assigned each portion of the dataset to select a candidate for code-switched question. Following this initial annotation, the annotators cross-checked each other’s work to identify any discrepancies. If a selected candidate appeared to be incorrect or suboptimal, the annotators engaged in thorough discussions until they reached an agreement on the most appropriate candidate. This iterative and collaborative process was integral to constructing a high-quality dataset.
@@ A.6 Dataset Size and Quality
KO
다음 글의 (가)에 대한 (나)의 상대적 특성으로 옳은 것은? (단, (가), (나)는 $각$ 겨울과 여름 중 하나임.)
우리나라는 더위와 추위에 대비하여 대청마루와 온돌 같은 전통 가옥 시설이 발달하였다. 대청마루는 바람을 잘 통하게 하여 (가) 을 시원하게 지낼 수 있도록 설치되었다. 온돌은 아궁이의 열을 방으로 전달하여 (나) 을 따뜻하게 지낼 수 있도록 설치되었다. 대청마루 는 중부와 남부 지역에 발달한 한편, 온돌은 대부분의 지역에 발달하였다.
(1) 평균 상대 습도가 높다.
(2) 정오의 태양 고도가 높다.
(3) 한파의 발생 일수가 많다.
(4) 대류성 강수가 자주 발생한다.
(5) 열대 저기압의 통과 횟수가 많다.
EN
What is the correct relative characteristic of (나) in relation to (가) in the following passage? (Note that (가) and (나) refer to either winter or summer.)
In Korea, traditional house facilities such as daecheongmaru and ondol have developed to cope with heat and cold. Daecheongmaru is designed to allow good ventilation to keep (가) cool. Ondol transfers heat from the kitchen stove to the room to keep (나) warm. While daecheongmaru is developed in the central and southern regions, ondol is developed in most areas.
(1) The average relative humidity is high.
(2) The midday sun’s altitude is high.
(3) There are many days of occurrence of cold waves.
(4) Heavy rainfall often occurs in Daeryuseong.
(5) There are many occurrences of passage of tropical cyclones.
CS
What is the correct relative characteristic of (나) in relation to (가) in the following passage? (Note that (가) and (나) refer to either winter or summer.)
In 한국, 전통 가옥 시설 such as 대청마루 and 온돌 have developed to cope with heat and cold. 대청마루 is designed to allow good ventilation to keep (가) cool. 온돌 transfers heat from the kitchen stove to the room to keep (나) warm. While 대청마루 is developed in the 중부 and 남부 지역, 온돌 is developed in most areas.
(1) The average 상대 습도 is high.
(2) The 정오의 태양 고도 is high.
(3) There are many days of occurrence of 한파.
(4) 대류성 강수 often occurs.
(5) There are many occurrences of passage of 열대 저기압.
@@ A.7 Data Sample
A.8 Resource Scarcity of Korean Language
There are varying perspectives on whether or not Korean is a low-resource language. The bottom line is, there is no clear consensus on the term ‘low-resource’ itself. Low-resource and high-resource are generally distinguished based on the availability of data resources and the size of the language’s speaker population (Nigatu et al., 2024). Some papers have classified Korean as low-resource (Lee et al., 2024), while others have described it as mid-resource (Joshi et al., 2020). What is certain is that Korean has fewer resources compared to high-resource languages (e.g., English, Chinese, Spanish, or Arabic), and particularly in the field of code-switching, there is only a handful of research or data. In our study, we aimed to compare cases where the LLM operates in the dominant language with cases where code-switching involving the target language is used. It is in this context that we referred to Korean as a low-resource language.
@@ C.2 Human Evaluator Qualifications
For knowledge identification evaluation, collecting qualified bilingual evaluators was not easy due to the inherent challenge in code-switching research of necessitating fluent bilinguals as evaluators. We managed to collect four Korean graduate school students as our evaluators, all of whom are native Korean with sufficient understanding of Korean culture. Also, they possess qualified English exam scores, indicating that they have no problem in understanding Korean-English code-switched texts. To mitigate the shortage of labor force, we designed the evaluation criteria objectively, allowing for an assessment that is not subjective and has clear correct answers. Specifically, we evaluate knowledge identification based on two criteria: faithfulness and helpfulness. Faithfulness evaluates the factualness of the knowledge, so the evaluators are required to use their background knowledge as well as searching from faithful sources where gold knowledge exists. To evaluate helpfulness, evaluators are given a gold answer to the question and determine whether the knowledge is helpful for finding the answer, using their logical reasoning.